# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 3.1 Two fixed points
Let ${\sf k}$ still be a field. We will need results from Graham–Kumar [25] on the $T$-equivariant ${\sf K}$-theory ${\sf K}_{T}(G/B)$ of $G/B$. Their $B$ is our $B^{+}$. As both $B$ and $B^{+}$ will be needed, let us consider the $G$-variety $\mathcal{B}$ of Borel subgroups. If $x$ is a $T$ fixed point in $\mathcal{B}$, then its stabilizer $B(x)$ is a Borel subgroup and one identifies $\mathcal{B}$ with $G/B(x)$. Say $x_{+}$ is the $T$ fixed point so that $B(x_{+})=B^{+}$. And $x_{-}$ is the fixed point so that $B(x_{-})=B$. Choose a representative $\dot{w_{0}}$ of $w_{0}$. One has $x_{+}=\dot{w_{0}}x_{-}$. We simply write $x_{+}={w_{0}}x_{-}$. Let $\phi_{0}$ be the isomorphism $B^{+}\to B$ sending $b$ to $\dot{w_{0}}b\dot{w_{0}}^{-1}$. If ${\mathcal{F}}$ is a $G$ equivariant vector bundle on $\mathcal{B}$, and $x$ is fixed by $T$, then the fibre ${\mathcal{F}}_{x}$ is a $B(x)$-module. We have ${\mathcal{F}}_{x_{+}}\cong\phi_{0}^{*}({\mathcal{F}}_{x_{-}})$. Recall that when $M$ is a finite dimensional $B$-module, we denote as in [29] by $\mathcal{L}(M)$ the $G$ equivariant vector bundle ${\mathcal{F}}$ with ${\mathcal{F}}_{x_{-}}=M$. If $M$ is a finite dimensional $B^{+}$-module, we denote by $\mathcal{L}^{+}(M)$ the $G$ equivariant vector bundle ${\mathcal{F}}$ with ${\mathcal{F}}_{x_{+}}=M$.
If ${\mathcal{F}}$ is a $T$ equivariant coherent sheaf on $G/B$ and $N$ is a finite dimensional $T$-module, then
@@
is a $T$ equivariant coherent sheaf.
When working with $B^{+}$ one should define excellent filtrations and relative Schubert filtrations in terms of $P^{+}(\lambda):=\phi_{0}^{*}(P(w_{0}\lambda))$ and $Q^{+}(\lambda):=\phi_{0}^{*}(Q(w_{0}\lambda))$. They have $B^{+}$ socles of weight $\lambda$ and are the costandard modules of [46].
@@
$H^{i}(B^{+},P^{+}(\lambda)\otimes_{\sf k}Q^{+}(\mu))=\begin{cases}{\sf k}&\text{if $i=0$ and $\lambda+\mu=0$}\\ 0&\text{else.}\end{cases}$
@@ 3.2 The Steinberg basis
The Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ consists of the corresponding one dimensional $B$-modules.
It provides by [44] a basis of the representation ring $R(T)$ as a module over the representation ring $R(G)={\sf K}_{0}({\mathrm{rep}}(G))$, where ${\mathrm{rep}}(G)$ is the category of finitely generated $G$-modules. We will find later that one still gets a basis if one replaces a few ${\sf k}_{e_{v}}$ by $Q(e_{v})$ or $P(-e_{v})^{*}$.
@@ 3.3 Schubert varieties and opposite Schubert varieties
Let the Schubert variety $X^{+}_{w}$ be the closure of $B^{+}wx_{+}$ and let the opposite Schubert variety $X^{w}$ be the closure of $B^{-}wx_{+}$. Its ‘boundary’ $\partial X^{w}$ is the union of the $X^{v}$ that are strictly contained in $X^{w}$.
The $T$-equivariant ${\sf K}$-theory ${\sf K}_{T}(G/B)$ of $G/B$ is a module for the representation ring $R(T)$. Generators of $R(T)$ are often written $e^{\lambda}$ instead of $[{\sf k}_{\lambda}]$. If $M$ is a finitely generated $T$-module, then its class $[M]$ in $R(T)$ is also written ${\mathrm{Char}}(M)$, the formal character of $M$.
If ${\mathcal{F}}$ is a $T$ equivariant coherent sheaf and $N$ is a finite dimensional $T$-module, then ${\mathcal{F}}\otimes_{\sf k}N$ represents $[N]\cdot[{\mathcal{F}}]$.
If ${\sf k}={\mathbb{C}}$, then we learn from [25] that the $R(T)$-module ${\sf K}_{T}(G/B)$ has a Schubert basis $\{[\mathcal{O}_{X^{+}_{w}}]\}_{w\in W}$. And also an ‘opposite Bruhat cell’ basis $\{[\mathcal{O}_{X^{w}}(-\partial X^{w})]\}_{w\in W}$. They are orthogonal under the $R(T)$-bilinear symmetric pairing $\langle$-,$\rangle$ - , - ⟩ on ${\sf K}_{T}(G/B)$ given by
$\langle[{\mathcal{F}}],[{\mathcal{G}}]\rangle=\sum_{i}(-1)^{i}{\mathrm{Char}}(H^{i}(G/B,{\mathcal{F}}\otimes{\mathcal{G}}))\in R(T).$
@@
If ${\mathcal{F}}\otimes{\mathcal{G}}$ is supported on a $T$ stable closed subscheme $Y$, then
$\langle[{\mathcal{F}}],[{\mathcal{G}}]\rangle=\sum_{i}(-1)^{i}{\mathrm{Char}}(H^{i}(Y,{\mathcal{F}}\otimes{\mathcal{G}}))\in R(T).$
@@
They are $G$-equivariant vector bundles, but we also view them as $T$-equivariant vector bundles.
@@
Before proving this, we need to recall things about Weyl groups.
4 Rappels: Weyl groups
4.1 The Demazure product
Let $(W,S=\{s_{\alpha_{1}},\dots,s_{\alpha_{r}}\})$ be the Coxeter presentation of the Weyl group of a simple simply connected group $G$ of rank $r$. The Demazure product on $W$ is defined by replacing the relation $s_{\alpha}\cdot s_{\alpha}=e$ for a simple root $\alpha$ with $s_{\alpha}\star s_{\alpha}=s_{\alpha}$. Equivalently, for any $w\in W$ and any simple root $\alpha$, define $w\star s_{\alpha}$ to be the longer of $ws_{\alpha}$ or $w$. A sequence of simple reflections is reduced if and only if the usual product $s_{\alpha_{i}}s_{\alpha_{i_{2}}}\dots s_{\alpha_{i_{k}}}$ is equal to the Demazure product $s_{\alpha_{i}}\star\dots\star s_{\alpha_{i_{k}}}$. If $v,w\in W$, then $\overline{BvB}~{}\overline{BwB}=\overline{Bv\star wB}$.
4.2 Total order $\prec$ on $W$, refining the Bruhat order
4.3 Combinatorics
First we recall some facts from [27] about $W$ and its action on $X(T)$. As usual $w_{0}$ denotes the longest element of $W$.
Lemma 4.1.
Let $v,w\in W$. The following are equivalent
  • $v<w$
  • $v^{-1}<w^{-1}$
  • $w_{0}w<w_{0}v$
  • $ww_{0}<vw_{0}$.
Proof.
For the first three, see [27, page 119]. Use that $(ww_{0})^{-1}=w_{0}w^{-1}$, $(vw_{0})^{-1}=w_{0}v^{-1}$. ∎
Lemma 4.2.
Proof.
See [27, page 116]. ∎
Lemma 4.3.
Proof.
See [27, page 16]. We usually prefer to multiply from the left. ∎
Lemma 4.4.
Let $w,z\in W$ with $\ell(wz)=\ell(w)+\ell(z)$. Then $w\leq w_{0}z^{-1}$.
Proof.
Lemma 4.5.
Every $w\in W$ may be written uniquely as $vz$, where $z\in W_{I}$ and $v\in W^{I}$ is the minimal coset representative of the coset $wW_{I}$.
Proof.
Compare [27, 1.15, page 22]. ∎
Recall that we denote the Demazure product on $W$ by $\star$.
Lemma 4.6.
Lemma 4.7.
Proof.
@@ 5 Triangularity of transition matrices
Recall that a weight $\lambda$ of a $B$-module $M$ is called extremal if all other weights $\mu$ of $M$ are shorter or equally long: $(\mu,\mu)\leq(\lambda,\lambda)$
@@
Let $S,S^{\prime}$ be unions of Schubert varieties $\overline{BuB}/B$ in $G/B$ and let $\lambda$ be dominant.
@@
  • $\Gamma(G/B,\mathcal{L}(\lambda))\to\Gamma(S,\mathcal{L}(\lambda))$ is surjective.
  • If $S^{\prime}\subset S$, then $\Gamma(S,\mathcal{L}(\lambda))\to\Gamma(S^{\prime},\mathcal{L}(\lambda))$ is surjective.
@@
See [47, Proposition 2.2.15, Lemma 2.2.11]. ∎
@@
We may assume ${\sf k}={\mathbb{C}}$, because of the base change properties [47, Chapter 7] of the $Q(\mu)$. Suppose $Q(-\mu)$ is larger than its socle ${\sf k}_{-\mu}$. Let $\nu$ be a weight of the socle of $Q(-\mu)/{\sf k}_{-\mu}$ and let $f$ be a nonzero weight vector of $Q(-\mu)$, of weight $\nu$, mapping to a vector of the socle of $Q(-\mu)/{\sf k}_{-\mu}$. As the socle of $Q(-\mu)$ is its weight space ${\sf k}_{-\mu}$, the vector $f$ can not be fixed by the unipotent radical $U$ of $B$. But it is well known that $U({\mathbb{C}})$ is generated by the $x_{-\alpha}(t)$ with $\alpha$ simple. So there must be such an $x_{-\alpha}(t)$ with $x_{-\alpha}(t)(f)-f$ a nonzero vector in ${\sf k}_{-\mu}$. In particular, $-\mu=\nu+n(-\alpha)$ for some integer $n$. But $\nu$ must be strictly shorter than the unique extremal weight $-\mu$ of $Q(-\mu)$. However, there is no strictly shorter weight in $-\mu+{\mathbb{Z}}\alpha$. ∎
@@
Let $w\in W$. If $M$ is a finite dimensional $B^{+}$-module with a relative Schubert filtration, then $H^{i}(\overline{B^{+}wx_{+}},\mathcal{L}^{+}(M))=0$ for $i>0$.
@@
    $-ve_{v}$ is an extremal weight of $\Gamma(\overline{Bw_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$, but not of $\Gamma((\partial\overline{BvB})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
@@
    If $-uve_{v}\neq-ve_{v}$, then $-uve_{v}$ is an extremal weight of $\Gamma((\partial\overline{BvB})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
@@
Part (1). The extremal weights of $\Gamma(\overline{Bw_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ are the elements in the $W$ orbit of $-w_{0}ve_{v}$, hence of $-ve_{v}$. Now suppose $-ve_{v}$ is an extremal weight of $\Gamma((\partial\overline{BvB})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$. Then $-ve_{v}$ can be written as $y(-w_{0}ve_{v})$ with $y\leq z\star v^{-1}w_{0}$ for some $z<v$. Replacing $z$ by a lesser element if necessary, we may assume $y\leq zv^{-1}w_{0}$ for some $z<v$. Put $u=yw_{0}$. Then $-ve_{v}=-uve_{v}$ with $u\geq zv^{-1}$. But $-ve_{v}=-uve_{v}$ implies $u\in W_{I}$, so $u\geq zv^{-1}$ implies $zv^{-1}\in W_{I}$. Thus $z^{-1}\in v^{-1}W_{I}$. This contradicts the minimality of $v^{-1}$ in its coset $v^{-1}W_{I}$.
@@
Let $v,w\in W$ such that $vw_{0}\nleq w$. Every extremal weight of $\Gamma(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ is also an extremal weight of $\Gamma((\partial\overline{Bww_{0}B})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
@@
Suppose $\ell((ww_{0})\star(v^{-1}w_{0}))=\ell(ww_{0})+\ell(v^{-1}w_{0})$. By Lemma 4.4 $(ww_{0})\leq v$, so $vw_{0}\leq w$, contrary to our assumption. So $\ell((ww_{0})\star(v^{-1}w_{0}))<\ell(ww_{0})+\ell(v^{-1}w_{0})$ and we have $y<ww_{0}$ with $(ww_{0})\star(v^{-1}w_{0}))=yv^{-1}w_{0}$, $\ell(yv^{-1}w_{0})=\ell(y)+\ell(v^{-1}w_{0})$. Every extremal weight of $\Gamma(\overline{Bww_{0}\star v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ is of the form $z(-w_{0}ve_{v})$ with $z\leq yv^{-1}w_{0}$. And $\Gamma((\partial\overline{Bww_{0}B})\;\overline{Bv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ maps onto $\Gamma(\overline{Byv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ and then onto $\Gamma(\overline{Bzx_{-}},\mathcal{L}(-w_{0}ve_{v}))$ by Lemma 5.1. ∎
Let $S$ be a union of Schubert varieties $\overline{BuB}/B$ in $G/B$ and let $w\in W$.
@@
Let ${\sf k}$ be a field of finite characteristic. Let $\hat{S}$ be the inverse image of $S$ in $G$.
@@
    The multiplication map $\overline{BwB}\times^{B}S\to G/B$ has connected fibres.
@@
    The multiplication map $\hat{S}\times^{B}\overline{BwB}/B\to G/B$ has connected fibres.

### v3

@@ 4.1 Two fixed points
Let ${\sf k}$ still be a field. We will need results from Graham–Kumar [28] on the $\mathbf{T}$-equivariant ${\sf K}$-theory ${\sf K}_{\mathbf{T}}(\mathbf{G}/\mathbf{B})$ of $\mathbf{G}/\mathbf{B}$. Their $\mathbf{B}$ is our $\mathbf{B}^{+}$. As both $\mathbf{B}$ and $\mathbf{B}^{+}$ will be needed, let us consider the $\mathbf{G}$-variety $\mathcal{B}$ of Borel subgroups. If $x$ is a $\mathbf{T}$ fixed point in $\mathcal{B}$, then its stabilizer $\mathbf{B}(x)$ is a Borel subgroup and one identifies $\mathcal{B}$ with $\mathbf{G}/\mathbf{B}(x)$. Let $x_{+}$ be the $\mathbf{T}$ fixed point the stabilizer of which is $\mathbf{B}(x_{+})=\mathbf{B}^{+}$. Similarly, let $x_{-}$ be the fixed point the stabilizer of which is $\mathbf{B}(x_{-})=\mathbf{B}$. Choose a representative $\dot{w_{0}}$ of $w_{0}$. One has $x_{+}=\dot{w_{0}}x_{-}$. We simply write $x_{+}={w_{0}}x_{-}$. Let $\phi_{0}$ be the isomorphism $\mathbf{B}^{+}\to\mathbf{B}$ sending $b$ to $\dot{w_{0}}b\dot{w_{0}}^{-1}$. If ${\mathcal{F}}$ is a $\mathbf{G}$-equivariant vector bundle on $\mathcal{B}$, and $x$ is fixed by $\mathbf{T}$, then the fibre ${\mathcal{F}}_{x}$ is a $\mathbf{B}(x)$-module. We have ${\mathcal{F}}_{x_{+}}\cong\phi_{0}^{*}({\mathcal{F}}_{x_{-}})$. Recall that when $M$ is a finite dimensional $\mathbf{B}$-module, we denote as in [32] by $\mathcal{L}(M)$ the $\mathbf{G}$-equivariant vector bundle ${\mathcal{F}}$ with ${\mathcal{F}}_{x_{-}}=M$. If $M$ is a finite dimensional $\mathbf{B}^{+}$-module, we denote by $\mathcal{L}^{+}(M)$ the $\mathbf{G}$-equivariant vector bundle ${\mathcal{F}}$ with ${\mathcal{F}}_{x_{+}}=M$.
If ${\mathcal{F}}$ is a $\mathbf{T}$-equivariant coherent sheaf on $\mathbf{G}/\mathbf{B}$ and $N$ is a finite dimensional $\mathbf{T}$-module, then
@@
is a $\mathbf{T}$-equivariant coherent sheaf.
When working with $\mathbf{B}^{+}$ one should define excellent filtrations and relative Schubert filtrations in terms of $P^{+}(\lambda):=\phi_{0}^{*}(P(w_{0}\lambda))$ and $Q^{+}(\lambda):=\phi_{0}^{*}(Q(w_{0}\lambda))$. They have $\mathbf{B}^{+}$-socles of weight $\lambda$ and are the costandard modules of [48].
@@
$H^{i}(\mathbf{B}^{+},P^{+}(\lambda)\otimes_{\sf k}Q^{+}(\mu))=\begin{cases}{\sf k}&\text{if $i=0$ and $\lambda+\mu=0$}\\ 0&\text{else.}\end{cases}$
@@ 4.2 The Steinberg basis
The Steinberg basis $\{{\sf k}_{e_{v}}\}_{v\in W}$ consists of the corresponding one dimensional $\mathbf{B}$-modules.
It provides by [46] a basis of the representation ring $R(\mathbf{T})$ as a module over the representation ring $R(\mathbf{G})={\sf K}_{0}({\mathrm{rep}}(\mathbf{G}))$, where ${\mathrm{rep}}(\mathbf{G})$ is the category of finitely generated $\mathbf{G}$-modules. We will find later (Corollary 8.25) that one still gets a basis if one replaces a few ${\sf k}_{e_{v}}$ by $Q(e_{v})$ or $P(-e_{v})^{*}$.
Generators of $R(\mathbf{T})$ are often written $e^{\lambda}$ instead of $[{\sf k}_{\lambda}]$. We will find later (Remark 10.3) that the Steinberg weights can also be described as follows: a weight $\lambda$ is a Steinberg weight if and only if $e^{\lambda}$ does not lie in the $R(\mathbf{G})$-submodule of $R(\mathbf{B})$ generated by the $e^{\mu}$ with $\mu<_{a}\lambda$. If $\lambda$ is a Steinberg weight, then $\lambda=e_{w}$ where $w\in W$ is the element of minimal length making $w\lambda$ dominant.
@@ 4.3 Schubert varieties and opposite Schubert varieties
Let the Schubert variety ${X}^{+}_{w}$ be the closure of $\mathbf{B}^{+}wx_{+}$ and let the opposite Schubert variety ${X}^{w}$ be the closure of $\mathbf{B}^{-}wx_{+}$. Its ‘boundary’ $\partial{X}^{w}$ is the union of the ${X}^{v}$ that are strictly contained in ${X}^{w}$.
The $\mathbf{T}$-equivariant ${\sf K}$-theory ${\sf K}_{\mathbf{T}}(\mathbf{G}/\mathbf{B})$ of $\mathbf{G}/\mathbf{B}$ is a module for the representation ring $R(\mathbf{T})$. If $M$ is a finitely generated $\mathbf{T}$-module, then its class $[M]$ in $R(\mathbf{T})$ is also written ${\mathrm{Char}}(M)$, the formal character of $M$.
If ${\mathcal{F}}$ is a $\mathbf{T}$-equivariant coherent sheaf and $N$ is a finite dimensional $\mathbf{T}$-module, then ${\mathcal{F}}\otimes_{\sf k}N$ represents $[N]\cdot[{\mathcal{F}}]$.
If ${\sf k}={\mathbb{C}}$, then we learn from [28] that the $R(\mathbf{T})$-module ${\sf K}_{\mathbf{T}}(\mathbf{G}/\mathbf{B})$ has a Schubert basis $\{[\mathcal{O}_{{X}^{+}_{w}}]\}_{w\in W}$. And also an ‘opposite Bruhat cell’ basis $\{[\mathcal{O}_{{X}^{w}}(-\partial{X}^{w})]\}_{w\in W}$. They are orthogonal under the $R(\mathbf{T})$-bilinear symmetric pairing $\langle$-,$\rangle$ - , - ⟩ on ${\sf K}_{\mathbf{T}}(\mathbf{G}/\mathbf{B})$ given by
$\langle[{\mathcal{F}}],[{\mathcal{G}}]\rangle=\sum_{i}(-1)^{i}{\mathrm{Char}}(H^{i}(\mathbf{G}/\mathbf{B},{\mathcal{F}}\otimes{\mathcal{G}}))\in R(\mathbf{T}).$
@@
If ${\mathcal{F}}\otimes{\mathcal{G}}$ is supported on a $\mathbf{T}$ stable closed subscheme $Y$, then
$\langle[{\mathcal{F}}],[{\mathcal{G}}]\rangle=\sum_{i}(-1)^{i}{\mathrm{Char}}(H^{i}(Y,{\mathcal{F}}\otimes{\mathcal{G}}))\in R(\mathbf{T}).$
@@
They are $\mathbf{G}$-equivariant vector bundles, but we also view them as $\mathbf{T}$-equivariant vector bundles.
@@ 5 Triangularity of transition matrices
Recall that a weight $\lambda$ of a $\mathbf{B}$-module $M$ is called extremal if all other weights $\mu$ of $M$ are shorter or equally long: $(\mu,\mu)\leq(\lambda,\lambda)$
A closed subset (i.e. reduced closed subscheme) of $\mathbf{G}/\mathbf{B}$ is $\mathbf{B}$-invariant if and only if it is a union of Schubert varieties. So there are only finitely many $\mathbf{B}$-invariant closed subsets of $\mathbf{G}/\mathbf{B}$.
@@
Let $S,S^{\prime}$ be unions of Schubert varieties in $\mathbf{G}/{\mathbf{B}}$ and let $\lambda$ be dominant.
@@
  • $\Gamma(\mathbf{G}/\mathbf{B},\mathcal{L}(\lambda))\to\Gamma(S,\mathcal{L}(\lambda))$ is surjective.
  • If $S^{\prime}\subset S$, then $\Gamma(S,\mathcal{L}(\lambda))\to\Gamma(S^{\prime},\mathcal{L}(\lambda))$ is surjective, and its kernel has a relative Schubert filtration.
  • $H^{i}(S,\mathcal{L}(\lambda))=0$ for $i>0$.
@@
See [49, Proposition 2.2.15, Lemma 2.3.10, Lemma 2.2.11, Proposition A.2.6], [10, Theorem 1.2.8, Chapter 2]. ∎
@@
We may assume ${\sf k}={\mathbb{C}}$, because of the base change properties [49, Chapter 7] of the $Q(\mu)$. Suppose $Q(-\mu)$ is larger than its socle ${\sf k}_{-\mu}$. Let $\nu$ be a weight of the socle of $Q(-\mu)/{\sf k}_{-\mu}$ and let $f$ be a nonzero weight vector of $Q(-\mu)$, of weight $\nu$, mapping to a vector of the socle of $Q(-\mu)/{\sf k}_{-\mu}$. As the socle of $Q(-\mu)$ is its weight space ${\sf k}_{-\mu}$, the vector $f$ can not be fixed by the unipotent radical $\mathbf{U}$ of $\mathbf{B}$. But it is well known that $\mathbf{U}({\mathbb{C}})$ is generated by the $x_{-\alpha}(t)$ with $\alpha$ simple. So there must be such an $x_{-\alpha}(t)$ with $x_{-\alpha}(t)(f)-f$ a nonzero vector in ${\sf k}_{-\mu}$. In particular, $-\mu=\nu+n(-\alpha)$ for some integer $n$. But $\nu$ must be strictly shorter than the unique extremal weight $-\mu$ of $Q(-\mu)$. However, there is no strictly shorter weight in $-\mu+{\mathbb{Z}}\alpha$. ∎
@@
Let $w\in W$. If $M$ is a finite dimensional $\mathbf{B}^{+}$-module with a relative Schubert filtration, then $H^{i}(\overline{\mathbf{B}^{+}wx_{+}},\mathcal{L}^{+}(M))=0$ for $i>0$.
@@
Notation 5.4.
If $v\in W$, then the boundary $\partial\overline{{\mathbf{B}}v\mathbf{B}}$ of $\overline{{\mathbf{B}}v\mathbf{B}}$ is defined to be the union of the $\overline{{\mathbf{B}}w\mathbf{B}}$ that are strictly contained in $\overline{{\mathbf{B}}v\mathbf{B}}$.
@@
    $-ve_{v}$ is an extremal weight of $\Gamma(\overline{{\mathbf{B}}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$, but not of $\Gamma((\partial\overline{{\mathbf{B}}v\mathbf{B}})\;\overline{{\mathbf{B}}v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
@@
    If $-uve_{v}\neq-ve_{v}$, then $-uve_{v}$ is an extremal weight of $\Gamma((\partial\overline{{\mathbf{B}}v{\mathbf{B}}})\;\overline{{\mathbf{B}}v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
@@
Part (1). The extremal weights of $\Gamma(\overline{{\mathbf{B}}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ are the elements in the $W$ orbit of $-w_{0}ve_{v}$, hence of $-ve_{v}$. Now suppose $-ve_{v}$ is an extremal weight of $\Gamma((\partial\overline{{\mathbf{B}}v{\mathbf{B}}})\;\overline{{\mathbf{B}}v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$. Then $-ve_{v}$ can be written as $y(-w_{0}ve_{v})$ with $y\leq z\star v^{-1}w_{0}$ for some $z<v$. Replacing $z$ by a lesser element if necessary, we may assume $y\leq zv^{-1}w_{0}$ for some $z<v$. Put $u=yw_{0}$. Then $-ve_{v}=-uve_{v}$ with $u\geq zv^{-1}$. But $-ve_{v}=-uve_{v}$ implies $u\in W_{I}$, so $u\geq zv^{-1}$ implies $zv^{-1}\in W_{I}$. Thus $z^{-1}\in v^{-1}W_{I}$. This contradicts the minimality of $v^{-1}$ in its coset $v^{-1}W_{I}$.
@@
Let $v,w\in W$ such that $vw_{0}\nleq w$. Every extremal weight of $\Gamma(\overline{{\mathbf{B}}(ww_{0}\star v^{-1}w_{0})x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ is also an extremal weight of $\Gamma((\partial\overline{{\mathbf{B}}ww_{0}{\mathbf{B}}})\;\overline{{\mathbf{B}}v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$.
@@
Suppose $\ell((ww_{0})\star(v^{-1}w_{0}))=\ell(ww_{0})+\ell(v^{-1}w_{0})$. By Lemma 3.6 $(ww_{0})\leq v$, so $vw_{0}\leq w$, contrary to our assumption. So $\ell((ww_{0})\star(v^{-1}w_{0}))<\ell(ww_{0})+\ell(v^{-1}w_{0})$ and we have $y<ww_{0}$ with $(ww_{0})\star(v^{-1}w_{0}))=yv^{-1}w_{0}$, $\ell(yv^{-1}w_{0})=\ell(y)+\ell(v^{-1}w_{0})$. Every extremal weight of $\Gamma(\overline{{\mathbf{B}}(ww_{0}\star v^{-1}w_{0})x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ is of the form $z(-w_{0}ve_{v})$ with $z\leq yv^{-1}w_{0}$. And $\Gamma((\partial\overline{{\mathbf{B}}ww_{0}{\mathbf{B}}})\;\overline{{\mathbf{B}}v^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ maps onto $\Gamma(\overline{{\mathbf{B}}yv^{-1}w_{0}x_{-}},\mathcal{L}(-w_{0}ve_{v}))$ and then onto $\Gamma(\overline{{\mathbf{B}}zx_{-}},\mathcal{L}(-w_{0}ve_{v}))$ by Lemma 5.1. ∎
Lemma 5.7.
Let $S^{\prime}$, $S^{\prime\prime}$ be unions of Schubert varieties with inverse images $\tilde{S}^{\prime}$, $\tilde{S}^{\prime\prime}$ respectively in $\mathbf{G}$. Let $w\in W$. Then $\tilde{S}^{\prime}(\overline{{\mathbf{B}}w{\mathbf{B}}}/{\mathbf{B}})\cap\tilde{S}^{\prime\prime}(\overline{{\mathbf{B}}w{\mathbf{B}}}/{\mathbf{B}})=(\tilde{S}^{\prime}\cap\tilde{S}^{\prime\prime})(\overline{{\mathbf{B}}w{\mathbf{B}}}/{\mathbf{B}})$.
Proof.
Clearly the right hand side is contained in the left hand side. Both sides are unions of Schubert varieties. First let $w$ be simple, say $w=s$. Consider $x\in W$ with $\overline{{\mathbf{B}}x{\mathbf{B}}}/{\mathbf{B}}$ contained in the left hand side. Choose $v$ minimal so that $v\star s=x\star s$. Use Lemma 3.9 to see that $\overline{{\mathbf{B}}(v\star s){\mathbf{B}}}/{\mathbf{B}}$ is contained in the right hand side. But then so is $\overline{{\mathbf{B}}x{\mathbf{B}}}/{\mathbf{B}}$. Next, if $\ell(w)>1$, choose a reduced expression $s_{1}\cdots s_{n}$ of $w$ and write $\overline{{\mathbf{B}}w{\mathbf{B}}}$ as $\overline{{\mathbf{B}}s_{1}{\mathbf{B}}}\cdots\overline{{\mathbf{B}}s_{n}{\mathbf{B}}}$. Show that $\tilde{S}^{\prime}(\overline{{\mathbf{B}}w{\mathbf{B}}})\cap\tilde{S}^{\prime\prime}(\overline{{\mathbf{B}}w{\mathbf{B}}})=(\tilde{S}^{\prime}\cap\tilde{S}^{\prime\prime})(\overline{{\mathbf{B}}w{\mathbf{B}}})$. ∎
Notation 5.8.
Let $S$ be a $B$-invariant closed subset union of $\mathbf{G}/{\mathbf{B}}$ and let ${\mathcal{F}}$ be a $\mathbf{B}$-equivariant vector bundle on $S$. If ${\sf k}$ is a field, put
$\chi(S,{\mathcal{F}})=\langle[\mathcal{O}_{S}],[{\mathcal{F}}]\rangle=\sum_{i}(-1)^{i}{\mathrm{Char}}(H^{i}(S,{\mathcal{F}}))$
in $R(T)$.
@@
Let $\lambda$ be a dominant weight and let $w\in W$. Let $\tilde{S}$ be the inverse image of $S$ in $\mathbf{G}$.
@@
    $H^{0}(\overline{{\mathbf{B}}w{\mathbf{B}}}S,\mathcal{L}(\lambda))\to H^{0}({\overline{{\mathbf{B}}w{\mathbf{B}}}/{\mathbf{B}}},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))$ is an isomorphism.
@@
    $H^{i}({\overline{{\mathbf{B}}w{\mathbf{B}}}/{\mathbf{B}}},\mathcal{L}(H^{0}(S,\mathcal{L}(\lambda)))=0$ for $i>0$.
  3. 3.
    If ${\sf k}$ is a field, then $\chi(\tilde{S}\overline{{\mathbf{B}}w{\mathbf{B}}}/\mathbf{B},\mathcal{L}(\lambda))=\chi(S,\mathcal{L}(H^{0}({\overline{{\mathbf{B}}w{\mathbf{B}}}/{\mathbf{B}}},\mathcal{L}(\lambda))).$